Ari Goldstein issued $700,000 of 15% 12 year bonds payable on january 1, 2016. The market interest rate at the date of issuance was 14% and the bonds pay interest semianually. Requirments

1. How much cash did the company receive upon issuance of bonds payable?

2. Prepare an amortization table for the bond using effective interest method, through the first two interest payments (round all numbers to the nearest whole dollar)

3. Journalize the issuance of the bonds on January 1 ,2016 and payments of the first semi annual interest amount and amortization of the bond on June 30, 2016 explanations are not required
Answer #1

Answer:

Part 1)

Cash received on issue=Present value of face value of bond+present value of coupon payments

=PVF(700000, n=12*2, r=14/2%)+PVAF(52500, n=12*2, r=14/2%)

Present Value (P)=C/{[(1+r)^n]}
Where,
C=Future Value= 700000
r=ROI= 0.07
n=No. of periods= 24
Now putting values in formula,
= 700000/{[(1+ 0.07)^24 ]}
= 700000/{[( 1.07)^24 ]}
= 700000/{[ 5.07237 ]}
= 138002.634
Present Value (P)=C/{[(1+r)^n]}
Where,
C=Future Value= 52500
r=ROI= 0.07
n=No. of periods= 24
Now putting values in formula,
= 52500/{[(1+ 0.07)^24 ]}
= 52500/{[( 1.07)^24 ]}
= 52500/{[ 0.19715 ]}
= 0.8029
= 52500{ 11.46933}
= 602140.03504

Now cash received=138,002.63+602,140.04=$740,142.67

Part 2)

E G A B с D F INTEREST INT EXPENSEL PAYMENT PREVIOUS (FACE BOOK AMORTIS BALANCE IN BOOK VAL 15/2 VAL IN ATION(B- PREMIUM(PR F

Part 3)

Date Account Titles and Explanation Jan 1, 16 Cash To Premium on issue of bonds To Bonds Payable Debit($) Credit($) 740,143 4

If helpful, Thumbs UP please:)